Dublin University Rifle Club

Version 1.1 Date : May 16, 2001

DURC Competition Attendance Report

Prone Indoor Rifle Nationals 2001

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Rathdrum Rifle and Pistol Club, Rathdrum
Date 22 April 2001
Discipline 25yd Prone Rifle

DURC Attendees

Mark Dennehy  Ray Reilly
John Keeney Conor MacDermottroe
Peter Byrne  Adrian Costigan
Joe Lawlor  Liam Murray
Alan Smyth  Deirdre Ninabar
Keith Butler  Ger Ryan

Scores

Mark Dennehy 587  Ray Reilly 580
John Keeney 574 Conor MacDermottroe 558
Peter Byrne 553  Adrian Costigan 548
Joe Lawlor 547  Liam Murray 540
Alan Smyth 537  Deirdre Ninabar 491
Keith Butler 484  Ger Ryan 469

Prizes Won

Mark Dennehy Gold Medal, National Championships

First Place, Class B

John Keeney First Place, Class C
Joe Lawlor First Place, Class D
Peter Byrne Second Place, Class D

Report

The Indoor Prone Rifle Nationals were held in the Rathdrum Rifle and Pistol Club range on April 22^{nd}. DURC attended with a large party of twelve competitors. Due to the large number of people attending (34 competitors from six clubs), the competition ran for nearly eight straight hours. Mark Dennehy shot a personal best, winning the Nationals with a lead of two points; this is the first time in the club’s history that we have taken the gold medal at this event, having had several silvers and bronze medals over the past twenty years. Mark also won the first place prize in Class B and was promoted on the spot to Class A. He is also one of the youngest people ever to win the Nationals. Mark is currently the DURC Secretary and serves on the National Target Shooting Association committee; he has been shooting for the past eight years with DURC. DURC may now count both the 25yd National Champion and the 50m National Champion amongst its members.

Also winning prizes were John Keeney, DURC Captain, who took first place in Class C with a score of 574, which will see him promoted to Class B for the next competition. Joe Lalor and Peter Byrne took first and second places in Class D, respectively; again, they face promotion before the next competition. In all, seven DURC shooters achieved scores high enough to see them promoted by a class before the next competition.

We also had several shooters attending their first ever external competition (Ger Ryan, Keith Butler, Deidre Ninaber, Liam Murray, Adrian Costigan and Joe Lalor) – that Joe was able to win his class outright on his first external competition says a great deal for his potential, as indeed do the scores the other newcomers produced on the day. We intend to promote their attendance at future external competitions.

Misé le meas,

Mark Dennehy, DURC Secretary